OpenStudy (anonymous):

The circular blade on a radial arm saw is turning at 277 rad/s at the instant the motor is turned off. In 22.0 s the speed of the blade is reduced to 85 rad/s. Assume the blade to be a uniform solid disk of radius 0.130 m and mass 0.400 kg. Find the net torque applied to the blade. Ugh?

OpenStudy (anonymous):

Torque=Moment of inertia * alpha alpha =change in omega /time=277-85/22 MI=mass*(radius)^2/2 multiply both to get the answer

OpenStudy (anonymous):

So 277-85/22=273.136=alpha (.4*.13^2)/2=.00338=MI .00338*273.136=.9232 .9232 is incorrect. What am I still doing wrong?

OpenStudy (jamesj):

\[ \alpha = \frac{\Delta \omega}{\Delta t} = \frac{277-85}{22} = \ ...\]

OpenStudy (jamesj):

(actually negative of that.)

OpenStudy (anonymous):

(277-85)/22=8.727 Torque=-.0295